A Clinical Psychologist wants to know what color is the most relaxing, to design new “calm”...

  1. A Clinical Psychologist wants to know what color is the most relaxing, to design new “calm” waiting rooms for dentists. Three different colors are viewed by participants and rated on a scale (0 – 10) to quantify the level of relaxation caused by the color. The n value and mean for each color are shown below, higher scores mean greater relaxation.

       Complete the source table below and answer the accompanying questions.  

                             Red                      Green                  Blue

                             n = 9                    n = 9                    n = 9

                             M = 1.9               M = 3.2               M = 4.3

Source                                   SS            df              MS        Fcalculated       Fcritical         p level         

Between Treatments     ___28__     ____       _____ ______ _____       p __ .05     

Within Treatments         __110_       ____       _____         

Total _______ _____

  1. What was the dependent variable ( D.V. ) in this study? __________________
  2. What was the independent variable (I.V.) in this study? ___________________
  3. Was there a significant effect? ______________________________________
  4. Write an APA formatted summary of this study below.
Answer #1

oral of S-1-4 = 64-4=60 2) dfxctment 2) af within afloral of Healment 2) Istal obserations = 64+1=65. 4) Foon ms treatment ms

please rate my answer and comment for doubts.